OpenStudy (anonymous):

i dont get it >_< someone help. solve the equation 2x - 3(-6x t 27) = 39

OpenStudy (anonymous):

Is that t supposed to be a +?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You need to expand by multiplying into the bracket.\[2x-3(-6x+27)=39\]\[2x+(-3)(-6x)+(-3)(27)=39\]\[2x+18x-81=39\]Can you continue from here by grouping like terms and solving for x?

OpenStudy (anonymous):

i know you have to do a number on both sides

OpenStudy (anonymous):

You need to bring all the terms with x in them to one side, and all those without an x to the other. Can you do that for me?

OpenStudy (anonymous):

oh i get it know ;) its 12.

OpenStudy (anonymous):

Nope! Try again. You subtracted 81 when you should have added it.

OpenStudy (anonymous):

Oh.!

OpenStudy (anonymous):

Or something.

OpenStudy (anonymous):

its 30.

OpenStudy (anonymous):

\[2x+18x-81=39\]Add the x terms together to get:\[20x-81=39\]Add 81 to both sides to get:\[20x=120\]Divide both sides by 20 to get:\[x=6\]